A jar contains 9 red marbles numbered 1 to 9 and 8 blue marbles numbered 1 to 8. A marble is drawn at random from the jar. Find the probability that the marble is blue AND even-numbered.



1
Expert's answer
2022-06-06T23:13:09-0400

We know that

"P(event)=\frac{number\: of\: favorable\: outcomes}{number\: of\: total\: outcomes}."


The total number of marbles is "9+8=17."

Among them we have 4 blue AND even-numbered marbles (numbered 2, 4, 6, 8).

So

P(getting blue AND even-numbered marble)"=\frac{4}{17}."


Answer: "\frac{4}{17}."